# Breaking the Visual Shortcuts in Multimodal Knowledge-Based Visual Question Answering

### v1

@@ 4.3 Document Encoder
Similarly to the query encoder, our objective is to obtain the document feature set $\mathcal{D}$. We assume that each document contains a text body and an image depicting the main entity, where the image is augmented following [deng-etal-2025-muka]. Given a collection of documents in KB, we further augment each document by collecting images of related entities mentioned within the document. We then extract sets of textual, visual, and multimodal features—analogous to the query encoding process—from both the document text body and the collected images, as detailed below.
Augmenting Documents with Images of Related Entities Given the textual content $T^{(d)}$ of a document, we first identify all named entities mentioned within $T^{(d)}$. For each extracted entity, we then collect the corresponding document from the KB, whose main entity matches that extracted entity. Since each document in the KB includes an image of its main entity, we extract these images and use them as related-entity images. By collecting these images, the augmented document becomes $(I^{(d)}_{0},I^{(d)}_{1},\dots,I^{(d)}_{R},T^{(d)})$, where $I^{(d)}_{0}$ is the image of the main entity and $I^{(d)}_{r}$ for $r\in\{1,\dots,R\}$ are the images of related entities. Note that while we rely on an LLM to identify related entities in order to maintain compatibility with the KB structure released in prior work [deng-etal-2025-muka], one could alternatively use the hyperlinks available within Wikipedia in a practical deployment.
@@
PreFLMR [lin2024preflmr] text 40.7 33.5 63.9 57.0 23.6 19.9 11.1 13.6
UniIR [wei2024uniir] text + single-image 34.5 26.4 22.0 25.7 13.8 14.6 4.7 11.2
ReT [caffagni2025recurrence] text + single-image 50.1 40.4 46.3 43.7 15.1 15.2 6.8 11.4
MuKA [deng-etal-2025-muka] text + single-image 51.7 41.0 64.9 60.7 20.8 18.7 8.6 12.2
@@ 5.1 Experimental Setup
Datasets We evaluate models on our RETINA dataset as well as two existing benchmarks, EVQA [mensink2023encyclopedic] and InfoSeek [chen2023can]. Since RETINA focuses on examples without visual shortcuts, whereas EVQA and InfoSeek include examples with visual shortcuts, the combination of these datasets enables a comprehensive assessment of realistic MKB-VQA scenarios. In terms of scale, RETINA contains 120K training samples and a 2K human-curated test set; EVQA provides 167K training samples and 3.7K test samples; and InfoSeek offers 676K training samples and 4.7K test samples. The knowledge bases of all three benchmarks consist of textual Wikipedia documents (details are available in Supp. Mat.) Following [deng-etal-2025-muka], we augment these documents with an image of their corresponding main entity. Building on this, we further enrich each document by attaching multiple images of related entities—identified as described in Sec. 4.3—thereby enabling multimodal retrieval with document embeddings augmented by multiple images.
Metrics We evaluate performance on both multimodal document retrieval and final answer generation. For retrieval, we adopt Recall@$K$, a standard metric for assessing retrieval systems, which measures the proportion of queries for which the annotated relevant document appears within the top-$K$ retrieved results. For answer generation, we use BERT Matching (BEM) [bulian-etal-2022-tomayto] on RETINA and EVQA, which scores a prediction by computing token-level matches weighted by contextual similarity from a BERT model, providing a soft semantic measure beyond strict lexical matching. For InfoSeek, we report the Accuracy metric specifically designed for this benchmark, which measures the proportion of correct answer predictions based on predefined matching functions tailored to each answer type [chen2023can]. All retrievers are evaluated by assessing the answers of the same VILA-13B [liu2023visualinstructiontuning] using their top-5 results following [deng-etal-2025-muka].
Models We compare our proposed MIMIR against four established multimodal baselines: UniIR [wei2024uniir], ReT [caffagni2025recurrence], PreFLMR [lin2024preflmr], and MuKA [deng-etal-2025-muka]. Among these, MIMIR, PreFLMR, and MuKA share largely the same core encoder architecture: they employ ColBERT-base-v2 for text encoding and CLIP ViT-G for visual encoding, along with their respective projection layers. The main difference lies in the design of the document encoder. PreFLMR operates on text-only documents, while MuKA enriches each document by attaching a single image corresponding to the main entity. In contrast, MIMIR further extends the document representation by incorporating multiple images of related entities within each document, as described in Sec. 4.3.
@@ 5.2 Results
Table 2: MIMIR Component Ablation. From top to bottom: MuKA [deng-etal-2025-muka]; extended MuKA with multi-image input; with multimodal features; and full MIMIR with entity token embeddings.
@@
Figure 6: Qualitative Comparison of Retrieval on the RETINA Benchmark. Results for (a) MuKA [deng-etal-2025-muka] and (b) MIMIR. GT documents are in green; distractors containing the query image entity are in red; query entity in GT documents are highlighted in green. While MuKA is biased toward visual shortcuts and often retrieves documents containing images similar to the query, MIMIR leverages multiple related image augmented document embeddings to correctly retrieve the GT document.
@@
Qualitative Analysis We analyze qualitative examples comparing MuKA [deng-etal-2025-muka] and MIMIR, examining how augmenting documents with multiple images of related entities affects performance on RETINA. MuKA tends to retrieve documents with images that are visually similar to the query, exhibiting a bias towards visual shortcut (e.g., trees or snowy landscapes in Fig. 6a). As a result, MuKA often fails to retrieve the GT document in RETINA, since the examples in our dataset are explicitly designed to eliminate such visual shortcuts. In contrast, MIMIR augments each document with multiple images of related entities, allowing the inclusion of an image that corresponds to the same entity depicted in the query, thereby enabling successful retrieval. As shown in Fig. 6b, related entities such as ‘Fraxinus americana’ and ‘Nunavut’ (marked in green) correspond to the entity of the query image, illustrating how this multi-image augmentation facilitates successful retrieval.
Further Analysis on Existing Benchmarks Interestingly, we observe consistent improvements over MuKA [deng-etal-2025-muka] on existing benchmarks that contain visual shortcuts, even though MuKA already augments each document with an image of its main entity. Our qualitative analysis reveals that even in benchmarks with visual shortcuts, the query image and the main-entity image in the GT document may differ visually (e.g., male vs. female wood ducks in Fig. 7a and 7b), which can lead to retrieval failures (e.g., MuKA results in Fig. 7c). In contrast, the proposed method further augments each document with images of related entities within the same GT document (e.g., entities highlighted in green in Fig. 7b), some of which are visually similar entities to the main entity (e.g., a mandarin duck in Fig. 7). This enrichment leads to a increased likelihood of retrieving the correct document. For example, Fig. 7d shows that the GT document for the wood duck is successfully retrieved because the document was augmented with the image of a mandarin duck—previously one of the most challenging negative samples for the baseline. This success is likely facilitated by structure of the Wikipedia, as related entities within it often depict visually similar entities.
@@ Appendix A RETINA Bench Construction
This section describes the process of constructing RETINA bench, including how textual queries and images were collected, processed, and curated to ensure quality. Throughout all stages of this pipeline, Qwen2.5-32B [bai2023qwen] was used.
@@ A.1 Query and Document Image Pool Construction
In this subsection, we explain how the pools of question and document images were created, including data sources, replacement strategies for unavailable images, and verification procedures to avoid duplication. Images for questions and documents were sourced from separate pools. Question images were taken from the M2KR dataset, including OVEN [hu2023opendomainvisualentityrecognition], Google Landmark v2 [weyand2020googlelandmarksdatasetv2], and iNaturalist [vanhorn2018inaturalistspeciesclassificationdetection]. Document images were downloaded from web page links provided by MuKA [deng-etal-2025-muka]. For images that were no longer available, replacements were retrieved using the Wikipedia API and Google Lens, and verified with the imagehash toolkit to ensure they were not identical to the query images, we used a perceptual hash size of 8, treating a distance of 10 or less as identical following [deng-etal-2025-muka].
@@ A.2 Paraphrasing
Here, we describe the paraphrasing process applied to the questions to prevent them from being directly retrievable using the text alone. We evaluate retrieval using BM25 [robertson1995okapi] with a top-5 setting. Before paraphrasing, approximately 37% of the noisy samples are retrievable. After paraphrasing, recall drops from 37% to 25%, and the remaining retrievable samples are removed from the dataset.
@@ Appendix B RETINA Qualitative Examples
PreFLMR [lin2024preflmr] text 40.7 63.9 23.6 11.1
UniIR [wei2024uniir] text + single-image 34.5 22.0 13.8 4.7
ReT [caffagni2025recurrence] text + single-image 50.1 46.3 15.1 6.8
MuKA [deng-etal-2025-muka] text + single-image 51.7 64.9 20.8 8.6
@@ Appendix C MIMIR Qualitative Examples
In this section, we illustrate qualitative examples of MuKA[deng-etal-2025-muka] (Fig. 11a) and MIMIR (Fig. 11b). Because RETINA is designed to avoid visual shortcuts, MuKA’s tendency to retrieve documents with images that are visually similar to the query leads it to select incorrect documents. In contrast, MIMIR augments the document with multiple images of related entities, allowing a visual match with the query. This visual evidence reinforces the textual relevance of the GT document, enabling successful retrieval.
@@ Appendix D KB Construction for Inference
In this section, we describe how the KB is constructed for retrieval, including the passages used and their configurations. EVQA and RETINA share the same KB, built from the EVQA subset of Wikipedia passages provided by the M2KR benchmark, while Infoseek uses a separate KB constructed from the Infoseek subset of M2KR Wikipedia passages [lin2024preflmr]. Since RETINA shares the same KB as EVQA, only samples whose GT documents do not appear in the EVQA training set are included as unseen test samples for RETINA. For KB indexing with PLAID [santhanam2022plaidefficientenginelate], we set the parameters as follows: the number of bits to 8, the number of k-means iterations to 20, used cosine similarity and the batch size was set to 64 following [santhanam2022colbertv2effectiveefficientretrieval, deng-etal-2025-muka, lin2024preflmr].
@@ Appendix E Implementation Details
In this section, we provide implementation details, including the dataset construction process and model configurations. Following MuKA [deng-etal-2025-muka], the model is initialized with the weights of PreFLMR [lin2024preflmr] using the CLIP ViT-G variant, which has demonstrated the best performance. Training was conducted for one epoch using the Adam optimizer. The projection layers are 2 layer MLP following [deng-etal-2025-muka, lin2024preflmr] for projecting the multimodal features as 32 tokens embeddings and lowering the dimension to 128 for efficiency [lin2023finegrainedlateinteractionmultimodalretrieval, lin2024preflmr, deng-etal-2025-muka]. For the number of augmented related images, we do not impose an upper bound on the number of augmented related images, which is an average 4.3 per sample. For training, we utilize 120k training sample randomly sampled from EVQA and RETINA, respectively. For the preliminary experiment, we discarded the text token features from the query embeddings while keeping all other architectural components and settings identical [deng-etal-2025-muka]. For answer generation, the VILA 13B model [lin2024vilapretrainingvisuallanguage] employed in [deng-etal-2025-muka] was used. The top-5 results from the zero-shot inference of the PreFLMR [lin2024preflmr] were utilized to perform LoRA training on VILA 13B. LoRA rank and alpha were set to 128 and 256, respectively, with a batch of 512. The Adam optimizer was applied, training only the multimodal projectors and LoRA modules with learning rates of 2e-5 and 2e-4 [deng-etal-2025-muka].

### v2

@@ 4.3 Document Encoder
Similarly to the query encoder, our objective is to obtain the document feature set $\mathcal{D}$. We assume that each document contains a text body and an image depicting the main entity, where the image is augmented following [6]. Given a collection of documents in KB, we further augment each document by collecting images of related entities mentioned within the document. We then extract sets of textual, visual, and multimodal features—analogous to the query encoding process—from both the document text body and the collected images, as detailed below.
Augmenting Documents with Images of Related Entities Given the textual content $T^{(d)}$ of a document, we first identify all named entities mentioned within $T^{(d)}$. For each extracted entity, we then collect the corresponding document from the KB, whose main entity matches that extracted entity. Since each document in the KB includes an image of its main entity, we extract these images and use them as related-entity images. By collecting these images, the augmented document becomes $(I^{(d)}_{0},I^{(d)}_{1},\dots,I^{(d)}_{R},T^{(d)})$, where $I^{(d)}_{0}$ is the image of the main entity and $I^{(d)}_{r}$ for $r\in\{1,\dots,R\}$ are the images of related entities. Note that while we rely on an LLM to identify related entities in order to maintain compatibility with the KB structure released in prior work [6], one could alternatively use the hyperlinks available within Wikipedia in a practical deployment.
@@
PreFLMR [19] text 40.7 33.5 63.9 57.0 23.6 19.9 11.1 13.6
UniIR [40] text + single-image 34.5 26.4 22.0 25.7 13.8 14.6 4.7 11.2
ReT [4] text + single-image 50.1 40.4 46.3 43.7 15.1 15.2 6.8 11.4
MuKA [6] text + single-image 51.7 41.0 64.9 60.7 20.8 18.7 8.6 12.2
@@ 5.1 Experimental Setup
Datasets We evaluate models on our RETINA dataset as well as two existing benchmarks, EVQA [27] and InfoSeek [5]. Since RETINA focuses on examples without visual shortcuts, whereas EVQA and InfoSeek include examples with visual shortcuts, the combination of these datasets enables a comprehensive assessment of realistic MKB-VQA scenarios. In terms of scale, RETINA contains 120K training samples and a 2K human-curated test set; EVQA provides 167K training samples and 3.7K test samples; and InfoSeek offers 676K training samples and 4.7K test samples. The knowledge bases of all three benchmarks consist of textual Wikipedia documents (details are available in Supp. Mat.) Following [6], we augment these documents with an image of their corresponding main entity. Building on this, we further enrich each document by attaching multiple images of related entities—identified as described in Sec. 4.3—thereby enabling multimodal retrieval with document embeddings augmented by multiple images.
Metrics We evaluate performance on both multimodal document retrieval and final answer generation. For retrieval, we adopt Recall@$K$, a standard metric for assessing retrieval systems, which measures the proportion of queries for which the annotated relevant document appears within the top-$K$ retrieved results. For answer generation, we use BERT Matching (BEM) [3] on RETINA and EVQA, which scores a prediction by computing token-level matches weighted by contextual similarity from a BERT model, providing a soft semantic measure beyond strict lexical matching. For InfoSeek, we report the Accuracy metric specifically designed for this benchmark, which measures the proportion of correct answer predictions based on predefined matching functions tailored to each answer type [5]. All retrievers are evaluated by assessing the answers of the same VILA-13B [21] using their top-5 results following [6].
Models We compare our proposed MIMIR against four established multimodal baselines: UniIR [40], ReT [4], PreFLMR [19], and MuKA [6]. Among these, MIMIR, PreFLMR, and MuKA share largely the same core encoder architecture: they employ ColBERT-base-v2 for text encoding and CLIP ViT-G for visual encoding, along with their respective projection layers. The main difference lies in the design of the document encoder. PreFLMR operates on text-only documents, while MuKA enriches each document by attaching a single image corresponding to the main entity. In contrast, MIMIR further extends the document representation by incorporating multiple images of related entities within each document, as described in Sec. 4.3.
@@ 5.2 Results
Table 2: MIMIR Component Ablation. From top to bottom: MuKA [6]; extended MuKA with multi-image input; with multimodal features; and full MIMIR with entity token embeddings.
@@
Figure 6: Qualitative Comparison of Retrieval on the RETINA Benchmark. Results for (a) MuKA [6] and (b) MIMIR. GT documents are in green; distractors containing the query image entity are in red; query entity in GT documents are highlighted in green. While MuKA is biased toward visual shortcuts and often retrieves documents containing images similar to the query, MIMIR leverages multiple related image augmented document embeddings to correctly retrieve the GT document.
@@
Qualitative Analysis We analyze qualitative examples comparing MuKA [6] and MIMIR, examining how augmenting documents with multiple images of related entities affects performance on RETINA. MuKA tends to retrieve documents with images that are visually similar to the query, exhibiting a bias towards visual shortcut (e.g., trees or snowy landscapes in Fig. 6a). As a result, MuKA often fails to retrieve the GT document in RETINA, since the examples in our dataset are explicitly designed to eliminate such visual shortcuts. In contrast, MIMIR augments each document with multiple images of related entities, allowing the inclusion of an image that corresponds to the same entity depicted in the query, thereby enabling successful retrieval. As shown in Fig. 6b, related entities such as ‘Fraxinus americana’ and ‘Nunavut’ (marked in green) correspond to the entity of the query image, illustrating how this multi-image augmentation facilitates successful retrieval.
Further Analysis on Existing Benchmarks Interestingly, we observe consistent improvements over MuKA [6] on existing benchmarks that contain visual shortcuts, even though MuKA already augments each document with an image of its main entity. Our qualitative analysis reveals that even in benchmarks with visual shortcuts, the query image and the main-entity image in the GT document may differ visually (e.g., male vs. female wood ducks in Fig. 7a and 7b), which can lead to retrieval failures (e.g., MuKA results in Fig. 7c). In contrast, the proposed method further augments each document with images of related entities within the same GT document (e.g., entities highlighted in green in Fig. 7b), some of which are visually similar entities to the main entity (e.g., a mandarin duck in Fig. 7). This enrichment leads to a increased likelihood of retrieving the correct document. For example, Fig. 7d shows that the GT document for the wood duck is successfully retrieved because the document was augmented with the image of a mandarin duck—previously one of the most challenging negative samples for the baseline. This success is likely facilitated by structure of the Wikipedia, as related entities within it often depict visually similar entities.
@@ Appendix A RETINA Bench Construction
This section describes the process of constructing RETINA bench, including how textual queries and images were collected, processed, and curated to ensure quality. Throughout all stages of this pipeline, Qwen2.5-32B [2] was used.
@@ A.1 Query and Document Image Pool Construction
In this subsection, we explain how the pools of question and document images were created, including data sources, replacement strategies for unavailable images, and verification procedures to avoid duplication. Images for questions and documents were sourced from separate pools. Question images were taken from the M2KR dataset, including OVEN [9], Google Landmark v2 [41], and iNaturalist [8]. Document images were downloaded from web page links provided by MuKA [6]. For images that were no longer available, replacements were retrieved using the Wikipedia API and Google Lens, and verified with the imagehash toolkit to ensure they were not identical to the query images, we used a perceptual hash size of 8, treating a distance of 10 or less as identical following [6].
@@ A.2 Paraphrasing
Here, we describe the paraphrasing process applied to the questions to prevent them from being directly retrievable using the text alone. We evaluate retrieval using BM25 [31] with a top-5 setting. Before paraphrasing, approximately 37% of the noisy samples are retrievable. After paraphrasing, recall drops from 37% to 25%, and the remaining retrievable samples are removed from the dataset.
@@ Appendix B RETINA Qualitative Examples
PreFLMR [19] text 40.7 63.9 23.6 11.1
UniIR [40] text + single-image 34.5 22.0 13.8 4.7
ReT [4] text + single-image 50.1 46.3 15.1 6.8
MuKA [6] text + single-image 51.7 64.9 20.8 8.6
@@ Appendix C MIMIR Qualitative Examples
In this section, we illustrate qualitative examples of MuKA[6] (Fig. 11a) and MIMIR (Fig. 11b). Because RETINA is designed to avoid visual shortcuts, MuKA’s tendency to retrieve documents with images that are visually similar to the query leads it to select incorrect documents. In contrast, MIMIR augments the document with multiple images of related entities, allowing a visual match with the query. This visual evidence reinforces the textual relevance of the GT document, enabling successful retrieval.
@@ Appendix D KB Construction for Inference
In this section, we describe how the KB is constructed for retrieval, including the passages used and their configurations. EVQA and RETINA share the same KB, built from the EVQA subset of Wikipedia passages provided by the M2KR benchmark, while Infoseek uses a separate KB constructed from the Infoseek subset of M2KR Wikipedia passages [19]. Since RETINA shares the same KB as EVQA, only samples whose GT documents do not appear in the EVQA training set are included as unseen test samples for RETINA. For KB indexing with PLAID [33], we set the parameters as follows: the number of bits to 8, the number of k-means iterations to 20, used cosine similarity and the batch size was set to 64 following [34, 6, 19].
@@ Appendix E Implementation Details
In this section, we provide implementation details, including the dataset construction process and model configurations. Following MuKA [6], the model is initialized with the weights of PreFLMR [19] using the CLIP ViT-G variant, which has demonstrated the best performance. Training was conducted for one epoch using the Adam optimizer. The projection layers are 2 layer MLP following [6, 19] for projecting the multimodal features as 32 tokens embeddings and lowering the dimension to 128 for efficiency [18, 19, 6]. For the number of augmented related images, we do not impose an upper bound on the number of augmented related images, which is an average 4.3 per sample. For training, we utilize 120k training sample randomly sampled from EVQA and RETINA, respectively. For the preliminary experiment, we discarded the text token features from the query embeddings while keeping all other architectural components and settings identical [6]. For answer generation, the VILA 13B model [15] employed in [6] was used. The top-5 results from the zero-shot inference of the PreFLMR [19] were utilized to perform LoRA training on VILA 13B. LoRA rank and alpha were set to 128 and 256, respectively, with a batch of 512. The Adam optimizer was applied, training only the multimodal projectors and LoRA modules with learning rates of 2e-5 and 2e-4 [6].